Suppose that Alicia has $30,000 in invested as follows: $10,000 in TJX Cos., $15,000 in Macy’s, and $5,000 in JC Penney. If the beta for these stocks is 0.84 for TJX, 1.06 for Macy’s, and 1.02 for JC Penney, the risk-free rate is 5% and the market risk premium is 8%, what is the expected return on Alicia’s portfolio?

(Please show work)

-Thank you!!
Answer #1

Portfolio of Beta =weight of TJX*Return of TJX +Weight of Macy's*Return of Macy's+Weight of JC Penney*Return of JC Penney
=10000/(10000+15000+5000)*0.84+15000/(10000+15000+5000)*1.06+5000/(10000+15000+5000)*1.02=0.98
Expected Return on Alicia's Portfolio =Risk free rate+Beta*Market Risk Premium =5%+0.98*8% =12.84%